# SegEarth-OV3: Exploring SAM 3 for Open-Vocabulary Semantic Segmentation in Remote Sensing Images

### v1

@@ 4 Experiments
| CLIP [48] | 17.7 | 3.5 | 19.6 | 16.0 | 7.7 | 3.9 | 4.9 | 7.1 | 18.6 |
| MaskCLIP [72] | 29.8 | 14.0 | 33.4 | 29.2 | 28.1 | 13.2 | 10.6 | 20.8 | 39.8 |
| SCLIP [55] | 33.4 | 21.0 | 34.9 | 25.9 | 21.1 | 7.0 | 7.4 | 14.9 | 32.1 |
| GEM [5] | 24.4 | 13.6 | 28.5 | 20.8 | 13.4 | 4.7 | 5.1 | 11.9 | 39.5 |
| ClearCLIP [33] | 36.6 | 20.8 | 39.0 | 30.1 | 25.5 | 5.7 | 6.4 | 16.3 | 44.9 |
| SegEarth-OV [40] | 49.2 | 28.4 | 44.6 | 37.0 | 35.4 | 17.8 | 11.5 | 23.8 | 60.2 |
@@ 4.1 Setup
Datasets. Following SegEarth-OV [40], we evaluate our method on 17 remote sensing datasets, covering diverse scenes, resolutions, and tasks.
We further test on nine binary extraction datasets focusing on critical geospatial objects: building extraction (WHUAerial [28], WHUSat.II [28], Inria [45], and xBD [25]), road extraction (CHN6-CUG [73], DeepGlobe [20], Massachusetts [46], and SpaceNet [54]), and flood detection (WBS-SI222https://www.kaggle.com/datasets/shirshmall/water-body-segmentation-in-satellite-images).
Additionally, we report results on general scene datasets (Pascal VOC20 [23], COCO Stuff [6], and Cityscapes [19]) to demonstrate universality. The evaluation metric is mIoU for multi-class segmentation and IoU of the foreground class for binary extraction tasks.
Implementation Details. We use the official SAM 3 model equipped with the Perception Encoder-Large+ (PE-L+) [4] backbone. Input images are resized to 10081008. Text prompts are derived directly from category names (e.g., “building”, “road”), and some categories are augmented with synonyms333The prompt setting is manually curated based on dataset features and has not been systematically explored in this version.. The background threshold and the initial confidence threshold for the Transformer decoder are manually tuned for each dataset to achieve roughly optimal performance. No test-time augmentation and extra post-processing method [1] is used.
4.2 Main Results
Semantic segmentation. We report the quantitative comparison on eight remote sensing semantic segmentation benchmarks in Table 4. Our method, SegEarth-OV3, achieves a new state-of-the-art, demonstrating a substantial performance leap over existing methods. Specifically, it achieves an average mIoU of 53.4%, surpassing the best previous training-free method, CorrCLIP (40.7% mIoU), by a remarkable margin of +12.7% mIoU. This performance also consistently outperformed training-based OVSS methods fine-tuned on remote sensing data, e.g., RSKT-Seg (39.7% mIoU) and Cat-Seg (39.1% mIoU). A notable observation is that our zero-shot method outperforms the fully supervised Oracle on some datasets. On UDD5 and VDD datasets, SegEarth-OV3 achieves 71.7% mIoU and 64.5% mIoU respectively, exceeding the Oracle baselines (56.5% and 62.9% mIoU). This surprising result suggests that the rich semantic knowledge and robust segmentation capabilities inherent in the SAM 3 foundation model can, in certain scenarios, exceed the generalization ability of models trained on domain-specific data. In Figure 4, we visualize the results on a large-scale remote sensing image.
@@ 4.2 Main Results
| TCL [12] | ViT-B/16 | 83.2 | 22.4 | 24.0 |
| CLIP-DINOiser [60] | 80.9 | 24.6 | 31.7 | |
| CoDe [59] | - | 23.9 | 28.9 | |
| CAT-Seg [18] | 94.6 | - | - | |
@@
| CLIP =[48] | ViT-B/16 | 41.9 | 4.4 | 5.0 |
| MaskCLIP [72] | 74.9 | 16.4 | 12.6 | |
| ClearCLIP [33] | 80.9 | 23.9 | 30.0 | |
| SCLIP [55] | 80.4 | 22.4 | 32.2 | |
| ProxyCLIP [34] | 80.3 | 26.5 | 38.1 | |
| LaVG [30] | 82.5 | 23.2 | 26.2 | |
| CLIPtrase [50] | 81.2 | 24.1 | - | |
| NACLIP [26] | 83.0 | 25.7 | 38.3 | |
| Trident [51] | 84.5 | 28.3 | 42.9 | |
| ResCLIP [67] | 86.0 | 24.7 | 35.9 | |
| SC-CLIP [2] | 84.3 | 26.6 | 41.0 | |
| CLIPer [52] | 85.2 | 27.5 | - | |
| CASS [31] | 87.8 | 26.7 | 39.4 | |
@@
| FreeDA [3] | ViT-L/14 | 87.9 | 28.8 | 36.7 |
| CaR [53] | 91.4 | - | - | |
| ProxyCLIP [34] | 83.2 | 25.6 | 40.1 | |
| ResCLIP [67] | 85.5 | 23.4 | 33.7 | |
| SC-CLIP [2] | 88.3 | 26.9 | 41.3 | |
| CLIPer [52] | 90.0 | 28.7 | - | |
@@ 4.3 Ablation Studies
4.4 Results on General Scene Datasets
@@ 5 Conclusion
This paper explores the potential of SAM 3 for remote sensing OVSS. Building upon the powerful zero-shot capabilities of the SAM 3 foundation model, we introduce an inference strategy to adapt its prompt-based architecture for remote sensing OVSS. By fusing semantic and instance outputs and leveraging presence-guided filtering, we effectively address the unique challenges of remote sensing. Our results on 20 diverse datasets show that SegEarth-OV3 not only sets a new training-free state-of-the-art but also outperforms supervised baselines in specific cases. These results highlight the significant potential of SAM 3 for specific domain tasks and demonstrate that, through appropriate adaptation strategies, training-free paradigms can serve as a powerful alternative to traditional supervised learning.

### v2

@@ 4 Experiments
| CLIP (Radford et al., 2021) | 17.7 | 3.5 | 19.6 | 16.0 | 7.7 | 3.9 | 4.9 | 7.1 | 18.6 |
| MaskCLIP (Zhou et al., 2022) | 29.8 | 14.0 | 33.4 | 29.2 | 28.1 | 13.2 | 10.6 | 20.8 | 39.8 |
| SCLIP (Wang et al., 2023) | 33.4 | 21.0 | 34.9 | 25.9 | 21.1 | 7.0 | 7.4 | 14.9 | 32.1 |
| GEM (Bousselham et al., 2024) | 24.4 | 13.6 | 28.5 | 20.8 | 13.4 | 4.7 | 5.1 | 11.9 | 39.5 |
| ClearCLIP (Lan et al., 2024) | 36.6 | 20.8 | 39.0 | 30.1 | 25.5 | 5.7 | 6.4 | 16.3 | 44.9 |
| SegEarth-OV (Li et al., 2025c) | 49.2 | 28.4 | 44.6 | 37.0 | 35.4 | 17.8 | 11.5 | 23.8 | 60.2 |
@@ 4.1 Setup
Datasets. We evaluate the performance of SegEarth-OV3 across three main open-vocabulary tasks, involving a total of 24 remote sensing benchmarks and 3 general image benchmarks to ensure a comprehensive assessment.
2D Remote Sensing Semantic Segmentation: We use 20 datasets in total. This includes eight multi-class remote sensing benchmarks (OpenEarthMap (Xia et al., 2023), LoveDA (Wang et al., 2021), iSAID (Waqas Zamir et al., 2019), Potsdam, Vaihingen111https://www.isprs.org/education/benchmarks/UrbanSemLab, UAVid (Lyu et al., 2020), UDD5 (Chen et al., 2018) and VDD (Cai et al., 2023)) and nine binary extraction datasets focusing on critical geospatial targets (buildings: WHUAerial (Ji et al., 2018), WHUSat.II (Ji et al., 2018), Inria (Maggiori et al., 2017), and xBD (Gupta et al., 2019); roads: CHN6-CUG (Zhu et al., 2021), DeepGlobe (Demir et al., 2018), Massachusetts (Mnih, 2013), and SpaceNet (Van Etten et al., 2018); and floods: WBS-SI222https://www.kaggle.com/datasets/shirshmall/water-body-segmentation-in-satellite-images). Specifically, we evaluate the performance of the OVSS method on three GF-series datasets, including the GID dataset (GF-2) (Tong et al., 2020), the GF-7 building dataset333https://figshare.com/articles/dataset/GF_LowGradeRoadDataset _zip/24457369/2?file=42966904, and the Low-Grade Road Dataset (GF-2) (Chen et al., 2023b).
3D Remote Sensing Semantic Segmentation: We utilize the STPLS3D (Chen et al., 2022b) dataset, performing multi-view projection from 2D predictions to 3D point clouds for evaluation.
The evaluation metric is mIoU for multi-class segmentation and IoU of the foreground class for binary segmentation tasks.
Implementation Details. We use the official SAM 3 model equipped with the Perception Encoder-Large+ (PE-L+) (Bolya et al., 2025) backbone. Input images are resized to 10081008. Text prompts are derived directly from category names (e.g., “building”, “road”), and some categories are augmented with synonyms. The background threshold and the initial confidence threshold for the Transformer decoder are slightly adjusted for each dataset to achieve roughly optimal performance. No test-time augmentation or extra post-processing method (Araslanov and Roth, 2020) is used.
| Method | GID | GF-7 Building | Low-Grade Road |
| CLIP | 22.7 | 19.1 | 3.3 |
| MaskCLIP | 35.4 | 26.1 | 9.5 |
| GEM | 40.1 | 23.3 | 4.9 |
| SCLIP | 37.6 | 27.2 | 5.7 |
| ClearCLIP | 39.7 | 28.2 | 5.1 |
| SegEarth-OV | 46.3 | 32.1 | 9.4 |
| ProxyCLIP | 42.8 | 34.9 | 22.1 |
| CorrCLIP | 40.9 | 32.2 | 10.5 |
| SegEarth-OV3 | 42.2 | 58.7 | 60.2 |
@@
| Method | Base Model | Comparator | WHU-CD | LEVIR-CD | S2Looking | |||
| PCA-KMeans (Celik, 2009) | / | KMeans | 14.3 | 7.7 | 10.0 | 5.2 | - | - |
| CVA (Bovolo and Bruzzone, 2006) | / | CVA Match | 7.2 | 3.7 | 9.3 | 4.9 | 5.8 | - |
| DCVA (Saha et al., 2019) | / | CVA Match | 20.6 | 11.5 | 13.9 | 7.4 | - | - |
| UCD-SCM (Tan et al., 2023) | FastSAM + CLIP | OTSU | 32.1 | 19.1 | 32.4 | 19.3 | - | - |
| AnyChange (Zheng et al., 2024) | SAM | Latent Match | 28.1 | 16.4 | 32.7 | 19.5 | 6.4 | - |
[…]
| Inst-CEG (Li et al., 2024a) | APE (Shen et al., 2024) | CEG | 62.5 | 45.5 | 63.3 | 46.3 | - | - |
[…]
| DynamicEarth (Li et al., 2026) | SAM2 + SegEarth-OV | DINOv2 | 57.4 | 40.2 | 50.5 | 33.8 | 37.6 | 23.1 |
| DynamicEarth (Li et al., 2026) | APE (Shen et al., 2024) | DINOv2 | 75.9 | 61.1 | 66.7 | 50.0 | 21.1 | 11.8 |
[…]
| SegEarth-OV3-CD (ours) | SegEarth-OV3 | JIPS (ours) | 88.0 | 78.6 | 85.3 | 74.3 | 48.8 | 32.3 |
| Training sets | Methods | mIoU (%) | Per Class IoU (%) | |||||
| Ground | Building | Tree | Car | Light pole | Fence | |||
| Real 3D Data | PointTransformer (Zhao et al., 2021) | 36.27 | 39.95 | 20.88 | 62.57 | 36.13 | 49.32 | 8.76 |
| RandLA-Net (Hu et al., 2020) | 42.33 | 46.13 | 24.23 | 72.46 | 53.37 | 44.82 | 12.95 | |
| SCF-Net (Fan et al., 2021) | 45.93 | 68.77 | 37.27 | 65.49 | 51.50 | 31.22 | 21.34 | |
| MinkowskiNet (Choy et al., 2019) | 46.52 | 64.22 | 29.95 | 61.33 | 45.96 | 65.25 | 12.43 | |
| KPConv (Thomas et al., 2019) | 45.22 | 60.87 | 32.13 | 69.05 | 53.80 | 52.08 | 3.40 | |
| Synthetic 3D Data | PointTransformer (Zhao et al., 2021) | 45.73 | 84.12 | 73.37 | 60.60 | 16.96 | 27.23 | 12.10 |
| RandLA-Net (Hu et al., 2020) | 45.03 | 76.78 | 57.74 | 56.08 | 28.44 | 40.36 | 10.78 | |
| SCF-Net (Fan et al., 2021) | 47.82 | 77.51 | 68.68 | 56.81 | 29.87 | 42.53 | 11.52 | |
| MinkowskiNet (Choy et al., 2019) | 50.78 | 85.23 | 72.66 | 64.80 | 31.31 | 36.85 | 13.83 | |
| KPConv (Thomas et al., 2019) | 49.16 | 85.50 | 70.65 | 63.84 | 28.75 | 32.97 | 13.22 | |
| Real+Synthetic 3D Data | PointTransformer (Zhao et al., 2021) | 47.64 | 80.19 | 76.35 | 57.13 | 36.35 | 23.72 | 12.10 |
| RandLA-Net (Hu et al., 2020) | 50.53 | 82.90 | 66.59 | 63.77 | 33.91 | 41.84 | 14.19 | |
| SCF-Net (Fan et al., 2021) | 50.65 | 77.80 | 58.98 | 64.86 | 46.37 | 40.50 | 15.41 | |
| MinkowskiNet (Choy et al., 2019) | 51.35 | 80.86 | 74.03 | 59.21 | 31.72 | 45.51 | 16.79 | |
| KPConv (Thomas et al., 2019) | 53.73 | 87.40 | 78.51 | 66.18 | 39.63 | 41.30 | 9.34 | |
| None | SegEarth-OV3 | 61.26 | 88.92 | 75.04 | 77.40 | 44.57 | 69.02 | 12.63 |
4.2 Main Results
Semantic segmentation. We report the quantitative comparison on eight remote sensing semantic segmentation benchmarks in Table 3.2. Our method, SegEarth-OV3, achieves a new state-of-the-art, demonstrating a substantial performance leap over existing methods. Specifically, it achieves an average mIoU of 53.4%, surpassing the best previous training-free method, CorrCLIP (40.7% mIoU), by a remarkable margin of +12.7% mIoU. This performance also consistently outperformed training-based OVSS methods fine-tuned on remote sensing data, e.g., RSKT-Seg (39.7% mIoU) and Cat-Seg (39.1% mIoU). A notable observation is that our zero-shot method outperforms the fully supervised Oracle on some datasets. On UDD5 and VDD datasets, SegEarth-OV3 achieves 71.7% mIoU and 64.5% mIoU respectively, exceeding the Oracle baselines (56.5% and 62.9% mIoU). We attribute this to the fact that fully supervised models are prone to overfitting when trained on domain-specific datasets with limited scale or variations. In contrast, this surprising result suggests that the rich semantic knowledge and robust segmentation capabilities inherent in the SAM 3 foundation model can provide superior generalization and robustness without requiring any target-domain supervision. In Figure 4, we visualize the inference results on a large-scale remote sensing image exceeding pixels. Crucially, this seamless large-scale segmentation prediction highlights the effectiveness of our presence-guided filtering in practical deployments. When performing sliding-window inference, querying a massive global vocabulary against small localized patches typically causes severe categorical hallucinations. By dynamically suppressing absent categories patch-by-patch, our filtering strategy effectively eliminates localized false positives, verifying its practical applicability in actual large-scale scenes.
@@ 4.2 Main Results
| TCL (Cha et al., 2023) | ViT-B/16 | 83.2 | 22.4 | 24.0 |
| CLIP-DINOiser (Wysoczańska et al., 2024) | 80.9 | 24.6 | 31.7 | |
| CoDe (Wu et al., 2024) | - | 23.9 | 28.9 | |
| CAT-Seg (Cho et al., 2024) | 94.6 | - | - | |
@@
| CLIP =Radford et al. (2021) | ViT-B/16 | 41.9 | 4.4 | 5.0 |
| MaskCLIP (Zhou et al., 2022) | 74.9 | 16.4 | 12.6 | |
| ClearCLIP (Lan et al., 2024) | 80.9 | 23.9 | 30.0 | |
| SCLIP (Wang et al., 2023) | 80.4 | 22.4 | 32.2 | |
| ProxyCLIP (Lan et al., 2025) | 80.3 | 26.5 | 38.1 | |
| LaVG (Kang and Cho, 2024) | 82.5 | 23.2 | 26.2 | |
| CLIPtrase (Shao et al., 2024) | 81.2 | 24.1 | - | |
| NACLIP (Hajimiri et al., 2025) | 83.0 | 25.7 | 38.3 | |
| Trident (Shi et al., 2025) | 84.5 | 28.3 | 42.9 | |
| ResCLIP (Yang et al., 2025) | 86.0 | 24.7 | 35.9 | |
| SC-CLIP (Bai et al., 2024) | 84.3 | 26.6 | 41.0 | |
| CLIPer (Sun et al., 2025) | 85.2 | 27.5 | - | |
| CASS (Kim et al., 2025) | 87.8 | 26.7 | 39.4 | |
@@
| FreeDA (Barsellotti et al., 2024) | ViT-L/14 | 87.9 | 28.8 | 36.7 |
| CaR (Sun et al., 2024) | 91.4 | - | - | |
| ProxyCLIP (Lan et al., 2025) | 83.2 | 25.6 | 40.1 | |
| ResCLIP (Yang et al., 2025) | 85.5 | 23.4 | 33.7 | |
| SC-CLIP (Bai et al., 2024) | 88.3 | 26.9 | 41.3 | |
| CLIPer (Sun et al., 2025) | 90.0 | 28.7 | - | |
@@ 4.3 Ablation Studies
4.4 Performance on Specialized Images
We evaluate SegEarth-OV3 on the GF-series datasets (Chen et al., 2022a) to test its performance on Chinese satellite images. As shown in Table 3, SegEarth-OV3 performs well on target-specific datasets. On the GF-7 Building and Low-Grade Road datasets, it achieves 58.7% and 60.2% IoU, respectively, outperforming previous CLIP-based methods (e.g., ProxyCLIP at 34.9% and 22.1%). However, on the GID dataset, the performance of SegEarth-OV3 (42.2% mIoU) is lower than that of SegEarth-OV (46.3% mIoU). The GID dataset consists of GF-2 satellite images, which have a relatively lower spatial resolution and distinct imaging characteristics compared to UAV or sub-meter commercial satellite images. This resolution difference introduces a notable domain gap for SAM 3, as the model is primarily trained on natural images with sharp object boundaries. Consequently, SAM 3 may struggle with the blurred textures and continuous macroscopic land-cover categories (e.g., forest and farmland) in GF-2 images. In contrast, CLIP-based methods like SegEarth-OV benefit from extensive pre-training on diverse web data, providing stronger semantic generalization for such low-resolution or continuous scenes. Nevertheless, for precise object extraction in higher-resolution images, SegEarth-OV3 remains a strong baseline.
4.5 Performance on OVCD Task
We evaluate the OVCD performance on the WHU-CD, LEVIR-CD, and S2Looking datasets. For the data processing, we crop the original test images of LEVIR-CD and WHU-CD into patches of pixels. For the S2Looking dataset, we use the original images directly for testing. Table 5 presents the quantitative comparison between our method and other change detection algorithms. Our method, which combines SegEarth-OV3 with the joint instance- and pixel-level verification strategy, achieves the highest performance across all three benchmarks. For example, it reaches 88.0% F1 and 78.6% IoU on WHU-CD, and 85.3% F1 and 74.3% IoU on LEVIR-CD. We observe two main findings from the results in Table 5. First, existing change detection architectures show consistent performance improvements when using SegEarth-OV3 as their segmentation backbone. For instance, the F1 score of DynamicEarth on WHU-CD increases from 75.9% to 79.7%, and on LEVIR-CD from 66.7% to 72.0% (Su et al., 2026). This demonstrates the high quality of the segmentation masks extracted by our SegEarth-OV3. Second, our proposed joint instance- and pixel-level verification strategy outperforms other matching algorithms (e.g., SFID and CACH) when using the same SegEarth-OV3 foundation. This confirms that combining instance-level topology with the pixel-level comparison effectively reduces the impact of recognition errors and spatial misalignments. It enables robust zero-shot change detection without requiring any temporal training.
4.6 Generalization to 3D Point Cloud Segmentation
To extend SegEarth-OV3 to 3D point cloud segmentation, we employ a multi-view projection strategy. Given a 3D point cloud and the corresponding camera intrinsic and extrinsic parameters, we project the 3D points onto multiple 2D image planes. We process these 2D images using SegEarth-OV3 to obtain semantic predictions, which are then projected back to the 3D points. Because a single 3D point may be visible in multiple overlapping camera views, the 2D predictions can sometimes be inconsistent due to occlusions or varied perspectives. To aggregate these predictions and determine the final semantic label for each point, we apply a straightforward class-weighted voting mechanism.
We evaluate this approach on the STPLS3D-WMSC dataset. As shown in Table 6, our training-free method achieves a 61.26% mIoU. It outperforms several fully supervised 3D models (e.g., PointTransformer, RandLA-Net, MinkowskiNet) trained on real and synthetic 3D data. This result demonstrates that the 2D visual representations from SegEarth-OV3 are highly robust across different viewpoints. By combining these 2D predictions with a simple weighted voting strategy, the model identifies 3D structures effectively without relying on 3D point cloud annotations or complex geometric refinements. In Figure 5, we provide a qualitative visualization of a 3D UAV scene. The result shows that our method accurately categorizes distinct targets in a complex terrain. The predicted point cloud closely matches the ground truth, confirming that the multi-view projection strategy successfully translates 2D semantic knowledge into 3D spatial topologies.
4.7 Performance on General Scene Datasets
@@ 5 Conclusion
This paper explores the application of SAM 3 for open-vocabulary remote sensing interpretation. We introduce SegEarth-OV3, a training-free framework that adapts the prompt-based architecture of SAM 3 for complex geospatial scenarios. For semantic segmentation, we fuse the semantic and instance outputs and apply presence-guided filtering to address the scale variations and category sparsity inherent in Earth observation images. Building upon these fused predictions, we extend the framework to OVCD. By introducing a joint instance- and pixel-level verification strategy, the method effectively mitigates errors caused by spatial misalignments and recognition inconsistencies across bi-temporal images. Furthermore, we demonstrate the generalization capability of SegEarth-OV3 by applying it to 3D point cloud segmentation via multi-view projection. Our comprehensive evaluation across 24 diverse remote sensing benchmarks shows that the proposed method establishes a strong training-free baseline for multi-dimensional remote sensing tasks. In certain scenarios, it even outperforms fully supervised models. These results confirm that, with appropriate adaptation mechanisms, general visual foundation models can serve as effective tools for open-vocabulary remote sensing analysis without requiring domain-specific training or complex architectures.